6. Propan Raya corporation’s outstanding bonds have $1,200 par value, a 8% Semiannual coupon, 8 years to maturity, and 8.5% YTM. What is the bond’s price? a) $ 1063.38 b) $876.00 c) $1165.68 d) $895.48 7. Banco Macro is expected to generate $150 million in free cash flow next year, and the free cash flow is expected to grow at a constant rate of 4% indefinitely. Banco Macro has no preferred stock or debt and its WACC is 6.85%. Branco Marco has 40 million in stock outstanding, what is its stock value per share? a)$125.00 b) $131.57 c)$13.16 d) $12.50 8. Pick the best opinion that corresponds with the impact of four actions (a, b, c, d) below on a bond’s YTM: a) Investors learn that the bonds are not subordinated to another debt issue____ b) The bond’s price decreases ________________ c) The bond is downgraded graded by the rating agencies________________ d) The economy seems to be shifting from a recession to boom____________ a). a, increases the YTM b. increases the YTM c. increases the YTM d. decreases the YTM b). a. decreases the YTM b. increases the YTM c. increases the YTM d. decreases the YTM c). a. decreases the YTM
Answer #1

1.
=1200*8%/8.5%*(1-1/1.0425^16)+1200/1.0425^16=1165.67907151798

2.
=150/(6.85%-4%)*1/40=131.578947368421

3.
b). a. decreases the YTM b. increases the YTM c. increases the YTM d. decreases the YTM